Project

General

Profile

RunPlan 22nd February 2017

Issues and tasks:

  • beam bump needs to be redone
  • EMR daq stability - looks fine
  • magnetic sweep - DONE
  • diffuser test

Wednesday 22, 16:00 - Thursday 23, 10:00

  • Tracker calibration
  • Reference run
  • Momentum scan with SSU on (206A), SSD on (206A), FC off:
    • Cycle through H17a, H17b, H17c, H17d
  • Beam line tuning postponed

Feb 20th to March 26th

Date Feb 20 Feb 21 Feb 22 Feb 23 Feb 24 Feb 25 Feb 26
MOM P. Franchini P. Franchini P. Franchini P. Franchini P. Franchini P. Franchini P. Franchini
BLOC Y.Song Y.Song Y.Song Y.Song Y.Song S. Boyd S. Boyd
SOC D. Rajaram D. Rajaram D. Rajaram D. Rajaram D. Rajaram D. Rajaram D. Rajaram
Lead Shifter (00:00-08:00) Y.Song Y.Song Y.Song Y.Song Y.Song E. Overton E. Overton
Shifter (00:00-08:00) C. Rogers M Courthold M Courthold E.Overton E.Overton
Lead Shifter(08:00-16:00) D. Rajaram D. Rajaram D. Rajaram D. Rajaram D. Rajaram S. Wilbur S.Wilbur
shifter 2 (08:00-16:00) H. Nebrensky H. Nebrensky H. Nebrensky H. Nebrensky H. Nebrensky M. Federova M.Federova
Lead shifter (16:00-00:00) A. Dobbs A. Dobbs A. Dobbs A. Dobbs A. Dobbs Y. Martyniak Y. Martyniak
Shifter 2 (16:00-00:00) K. Long K. Long C. Hunt K. Long K. Long M. Savic M. Savic
Additional Personnel
Notes

Data to take:

Field On Data

Task H17 - Momentum scan with SSU and SSD on and FC off

Task Momentum # Particle Triggers Beamline Tag CoolingChannel Tag Similar runs Final Event Tally (k)
H17a 140 630K TOF2 3-140+M3-Test2 2016-05-1-SSUSSD 290K
H17b 170 1.5e5 TOF1 3-170+M3-Test1 2016-05-1-SSUSSD 150K
H17c 200 1.5e5 TOF1 3-200+M3-Test1 2016-05-1-SSUSSD 196K
H17d 240 1.5e5 TOF1 3-240+M3-Test1 2016-05-1-SSUSSD 150K